OpenStudy (anonymous):

Help, test tomorrow! Derivatives and Chain Rule derivative of y=csc^-1(secx), 0<=x<(pi/2)

OpenStudy (anonymous):

\[y = \csc ^{-1}(secx) , 0<=x<\frac{ \pi }{ 2 }\]

OpenStudy (anonymous):

i tried to set it up, but somewhere i went wrong and now im lost

OpenStudy (anonymous):

Next I did, \[y \prime = -\frac{ 1 }{ \left| secx \right| \sqrt{\sec ^{2}x-1}}*secx*tanx*secx*tanx\]

OpenStudy (anonymous):

and then when simplifying, i got lost with \[y \prime = -\frac{ \sec ^{2}xtanx }{ secx \sqrt{\sec ^{2}x-1} }\]

OpenStudy (anonymous):

can someone help me?